The EM Algorithm and Extensions. Geoffrey J. McLachlan and Thriyambakam Krishnan. Ser. in Prob. & Stat. Wiley, 1997, xvii + 274 pp, $59.95 [ISBN 0-471-12358-7].

Excellent resource for theory and application of EM algorithm and its many variations. Applications range from simple, one-parameter multinomials, to hidden Markov models, epidemiological models, and neural networks. No exercises. MK